OpenStudy (anonymous):

I encountered these two types of arrows that my instructors have been drawing every time they start a new line, and sometimes they draw one and not the other. I know one stands for "this implies", which is "=>" The other stands for "if and only if", which is "<=>" However, what I am having trouble with is when I am doing a general math problem, how do I know if I am to use one in the line and not the other, or use both?

OpenStudy (akshay_budhkar):

depends on what you are solving mostly you will encounter the implies sign which is used to solve a question whose solution has a number of steps following in a correct sequence. the second one is used only if there is a condition in that step like say x=3 if and only if x+1=y or something like that

OpenStudy (anonymous):

"if and only if", or "iff" for short is a lot stronger than just a one directional arrow. To prove that :\[p\iff q\] You need to prove two things:\[p\Longrightarrow q\]and\[q\Longrightarrow p\] Not every statement is an "iff" statement. for example:\[x^2=1\iff x=1\]would be incorrect because it is true that:\[x=1\Longrightarrow x^2=1\]it is NOT true that:\[x^2=1\Longrightarrow x=1\]x could be -1 as well.